### v1

@@ 1 Introduction
Heavy-ion collisions recreate a hot and dense, deconfined state of nuclear matter known as the Quark-Gluon Plasma (QGP). In contrast, small systems like proton-proton, proton-nucleus, and -nucleus collisions are not expected to form a hot, thermalized medium. Nevertheless, some QGP-like signatures, such as the suppression of quarkonium, can occur due to medium-induced modifications. These phenomena are referred to as cold nuclear matter (CNM) effects.
Disentangling CNM effects from the QGP is crucial for accurately interpreting heavy-ion collision data. Additionally, CNM effects play a significant role in understanding nucleon and nucleus structure.
The CNM effects are quantified using the nuclear modification factor: , where is the Parton distribution function (PDF) of a parton flavour “” inside a proton bound in a nucleus, and is PDF of a free proton. A is the number of nucleon of the target nucleus. PDF depends on both Bjorken and momentum transfer scale . A value of indicates the absence of nuclear effect [1].
@@ 2 Physics of UPCs
In ultra-peripheral collisions (UPCs), ions interact via electromagnetic processes, with no hadronic interactions taking place. This occurs when two relativistic heavy ions interact with an impact parameter larger than the sum of the radii of the colliding nuclei. Experimentally, UPC events are characterized by only a few tracks in an otherwise empty detector, a feature that applies specifically to exclusive (double-gap) events. Examples include exclusive (Pb-Pb) and dissociative (p-Pb) vector meson photoproduction.
In the exclusive process, both the projectile and the target nuclei remain intact. This process can be further categorized into two sub-classes: coherent process (photon interacts with the entire nucleus) and incoherent (where the photon interacts with individual nucleons or groups of nucleons within the nucleus).
This article discusses photoproduction cross-section measurements as a function of rapidity (), transverse momentum (Mandelstam) and photon-nucleon center-of-mass energy (). These variables are interrelated and also connect to Bjorken as shown in equations 1 - 3.
@@ 4.1 Coherent measurement at forward and midrapidity
At midrapidity three decay channels of J/ and are explored thanks to excellent PID of the TPC. They are , and . Good agreement is found with theory predictions which include gluon shadowing effects of about 0.65 at for and . The b-BK model which includes gluon saturation effects qualitatively describes the data within uncertainty [13].
4.2 Energy dependence of photonuclear cross section
@@
Energy dependence of coherent photonuclear cross section and nuclear suppression factor are measured in UPC Pb-Pb collisions at = 5.02 TeV using the above mentioned approach [14]. As shown in Fig 1 (a)., results includes both mid and forward rapidity data. Measurement covers GeV with Bjorken interval . No model describe the data completely. STARLight [7] and the impulse approximation [6] describe well low energy data, at high energy data are described by both models that include shadowing (EPS09-LO [8], LTA [9]) and saturation effects (GG-HS)[10].
The measured nuclear suppression factor gives quantitative values of nuclear shadowing and are consistent with the previous ALICE measurements at midrapidity [12]. Fig 2 (a) shows that suppression gets stronger with increasing center of mass energy. Impulse approximation without any nuclear effect overshoots the data [14].
4.3 Energy dependence of exclusive measurement
@@
4.4 Mandelstam dependence of coherent measurement
To search for gluon saturation, -dependent measurements are effective observables. The impact parameter (b) and the transverse momentum of vector mesons are Fourier conjugates, making the -dependent photonuclear cross section sensitive to the spatial gluon distribution. The coherent cross section specifically probes the average gluon density [22, 23]. ALICE measurements at = 5.02 TeV in Pb-Pb UPCs reveal that STARLight overestimates the data, while models such as LTA (accounting for shadowing effects) and b-BK (incorporating gluon saturation effects) qualitatively agree with the results [25].
@@ 5 Incoherent and dissociative measurements with ALICE
As explained in Section 2, in incoherent processes, the photon interacts with individual nucleons, characterized by transverse momentum () values larger than 200 MeV/c. The average is inversely proportional to the size of the target: the entire nucleus for coherent interactions, nucleons for incoherent processes in Pb-Pb collisions, and protons in dissociative interactions for p-Pb collisions.
5.1 Mandelstam dependence of incoherent measurement
As discussed in Section 4.4, -dependent measurements are effective probes of gluon saturation. The -dependent incoherent photoproduction can be used to study the variance (quantum fluctuations) [24]. For heavy nuclei, saturation is expected at higher-, where the photon interacts with the sub-nucleonic structure, making sensitive to quantum fluctuations [22]. The incoherent J/ photonuclear cross section was measured as a function of Mandelstam in Pb-Pb collisions at = 5.02 TeV. Models incorporating quantum fluctuations of the gluon density describe the data better than those without, although no model fully reproduces the normalization from proton to nuclear targets. The slope of the -dependence incoherent cross section provides sensitivity to spatial gluon fluctuations, allowing for the first-time probing of gluonic “hot spots” in Pb nuclei [26].
5.2 Energy dependence of dissociative J/ measurement
In dissociative processes, the photon interacts with the subnucleonic structure, making sensitive to the variance rather than the average of the gluon distribution [21]. Dissociative production provides access to subnucleonic fluctuations inside the proton [22].
HERA data does not cover the full kinematic range accessible at the LHC due to lower energies. The first measurement at collider energies was performed in p-Pb collisions at = 8.16 TeV. ALICE results are consistent with H1 data for absolute cross sections and agree well with the CCT “hot spots” model, which predicts a maximum cross section at 500 GeV. These measurements serve as a probe for subnucleonic fluctuations inside the proton [19].
@@ 6 UPC physics prospects with ALICE experiment in Run 3 and beyond
ALICE has undergone a significant upgrade, including the installation of the Forward Interaction Trigger (FIT) to replace the V0, T0, and AD detectors, as well as new ITS2 and forward detectors like MID (replacing the Muon Trigger) and the Muon Forward Tracker (MFT). Other improvements include enhanced readouts for TOF, ZDC, and the MUON arm, a new Central Trigger Processor, and an upgraded DAQ/Offline system with continuous readout [4, 5]. The Forward Calorimeter (FoCal) detector, planned for Run 4 (starting in 2029), will be positioned 7 m from IP2 on the on the opposite side of the muon arm, covering [27].
In Run 3, more flexible UPC event selection is enabled with FIT and ZDC signals, allowing exclusive, inclusive, and semi-inclusive (jets and diffractive) event studies. The progression in statistics from Run 1 (80 J/ events) to Run 2 (22k events) and Run 3 (70k events) reflects a significant increase at forward rapidity [12, 28]. Projections in the CERN Yellow Report estimate 1.1M J/ at midrapidity and 600k at forward rapidity in Run 3 and Run 4 at LHC energies [29].
The increased integrated luminosity enables precision studies of vector meson photoproduction in UPCs, reducing uncertainties for nuclear suppression factors to [29], Fig 2 (a). New opportunities include double differential studies, double vector meson photoproduction, UPC bottomonia production, and dissociative J/ in Run 3 with FoCal acceptance in Run 4. Possibly, Run 3 will also allow exclusive J/ and studies in p-Pb UPCs, with further enhancements expected in Run 4 using FoCal acceptance as projected in this article [30].
7 Summary and outlook
ALICE has made significant measurements probing QCD effects at high energies through coherent, exclusive, and dissociative J/ production. These studies provide insights into gluon shadowing, gluon saturation, and subnucleonic fluctuations, offering valuable comparisons with theoretical models. The ongoing Run 3 and upcoming Run 4 present exciting opportunities, with new data enabling precision studies and novel measurements, such as inclusive open charm and J/ production in UPCs.

### v2

@@ 1 Introduction
Heavy-ion collisions recreate a hot and dense, deconfined state of nuclear matter known as the Quark-Gluon Plasma (QGP). In contrast, small systems like proton-proton, proton-nucleus, and -nucleus collisions are not expected to form a hot, thermalized medium. Nevertheless, some QGP-like signatures, such as the suppression of quarkonium, can occur due to medium-induced modifications. These phenomena are referred to as cold nuclear matter (CNM) effects. Disentangling CNM effects from the QGP is crucial for accurately interpreting heavy-ion collision data. Additionally, CNM effects play a significant role in understanding nucleon and nucleus structure.
The CNM effects are quantified using the nuclear modification factor: , where is the Parton Distribution Function (PDF) of a parton flavour “” inside a proton bound in a nucleus, and is the PDF of a free proton. A is the number of nucleons of the target nucleus. The PDF depends on both Bjorken and momentum transfer scale . A value of indicates the absence of nuclear effect [1].
@@ 2 Physics of UPCs
In ultra-peripheral collisions (UPCs), ions interact via electromagnetic processes, with no hadronic interactions taking place. This occurs when two relativistic heavy ions interact with an impact parameter larger than the sum of the radii of the colliding nuclei. Experimentally, UPC events are characterized by only a few tracks in an otherwise empty detector, a feature that applies specifically to exclusive (double-gap) events. Examples include exclusive (Pb-Pb) and dissociative (p-Pb) vector meson photoproduction. In the exclusive process, both the projectile and the target nuclei remain intact. This process can be further categorized into two sub-classes: coherent process (photon interacts with the entire nucleus) and incoherent (where the photon interacts with individual nucleons or groups of nucleons within the nucleus).
This article discusses photoproduction cross-section measurements as a function of rapidity (), transverse momentum (Mandelstam) and photon-nucleon center-of-mass energy (). These variables are interrelated and also connect to Bjorken as shown in equations 1 - 2.
@@ 4 Coherent and exclusive measurements with ALICE
Energy dependence of coherent photonuclear cross section and nuclear suppression factor are measured in UPC Pb-Pb collisions at = 5.02 TeV using the above mentioned approach [13]. As shown in Fig 1 (a), results includes both mid and forward rapidity data. The data covers GeV with Bjorken interval . No model describe the data completely. STARLight [7] and the impulse approximation [6] describe well low energy data, at high energy data are described by both models that include shadowing (EPS09-LO [8], LTA [9]) and saturation effects (GG-HS)[10].
@@ 5 Incoherent and dissociative measurements with ALICE
To search for gluon saturation, -dependent measurements are effective observables. The -dependent incoherent photoproduction can be used to study the variance (quantum fluctuations). For heavy nuclei, saturation is expected at higher-, where the photon interacts with the sub-nucleonic structure, making sensitive to quantum fluctuations [21]. The incoherent J/ photonuclear cross section was measured as a function of Mandelstam in Pb-Pb collisions at = 5.02 TeV. Models incorporating quantum fluctuations of the gluon density describe the data better than those without, although no model fully reproduces the normalization from proton to nuclear targets. The slope of the -dependence incoherent cross section provides sensitivity to spatial gluon fluctuations, allowing for the first-time probing of gluonic “hot spots” in Pb nuclei [22].
In dissociative processes, the photon interacts with the subnucleonic structure, making sensitive to the variance rather than the average of the gluon distribution [20]. Dissociative production provides access to subnucleonic fluctuations inside the proton [21]. HERA data does not cover the full kinematic range accessible at the LHC due to lower energies. The first measurement at collider energies was performed in p-Pb collisions at = 8.16 TeV. ALICE results are consistent with H1 data for absolute cross sections and agree well with the CCT “hot spots” model, which predicts a maximum cross section at 500 GeV. These measurements serve as a probe for subnucleonic fluctuations inside the proton [18].
@@ 6 UPC physics prospects with ALICE experiment in Run 3 and beyond
ALICE has undergone a significant upgrade, including new detectors, enhanced readouts, a new Central Trigger Processor, and an upgraded DAQ/Offline system [4, 5]. The Forward Calorimeter (FoCal) detector, planned for Run 4 (starting in 2029), will be positioned 7 m from IP2 on the on the opposite side of the muon arm, covering [23].
In Run 3 and 4, the increased integrated luminosity enables precision studies of vector meson photoproduction in UPCs, reducing uncertainties for nuclear suppression factors to [24], Fig 2 (a). New opportunities include double differential studies, double vector meson photoproduction, UPC bottomonia production, and dissociative J/ in Run 3 with FoCal acceptance in Run 4. Run 3 might also allow exclusive J/ and studies in p-Pb UPCs, with further enhancements expected in Run 4 using FoCal acceptance as projected in this article [25].